OpenStudy (anonymous):

what is the point slope equation of the line that has a slope of -5 and goes tru the point (-4,9)?

OpenStudy (anonymous):

Which two options could be the next step in simplifying this expression? 27 + (3 + 41) A. 27 + 44 (use order of operations) B. 30 + 68 (use the distributive property) C. 30 + 41 (use the associative property) D. 68 + 30 (use the commutative property)

OpenStudy (anonymous):

point slope equation is y- y1=m(x-x1) m = slope or -5. the y1 is your y value in the point (-4,9) y1=9 x1 = -4 ( the x value in the point ) When you plug in the value it should look like this: \[y-9=-5(x-(-4) )\] Two negatives equal a positive so your final answer is: \[y-9=-5(x+4 )\]

OpenStudy (anonymous):

Which two options could be the next step in simplifying this expression? 27 + (3 + 41) A. 27 + 44 (use order of operations) B. 30 + 68 (use the distributive property) C. 30 + 41 (use the associative property) D. 68 + 30 (use the commutative property)

OpenStudy (anonymous):

but it says math procesing error